OpenStudy (gelow):

What is the solution to the equation 3/4 (4c + 16) = 2c +9 ? c =

OpenStudy (gelow):

@CloverRacer @AndrewTheCookie @chosenmatt @KamiBug @lizz123 @TheSmartOne

OpenStudy (gelow):

@ayyookyndall

OpenStudy (lizz123):

http://www.geteasysolution.com/3/4(4c+16)=2c+9 3/4(4c+16)=2c+9 We simplify the equation to the form, which is simple to understand 3/4(4c+16)=2c+9 Reorder the terms in parentheses +(+3c+12)=2c+9 Remove unnecessary parentheses +3c+12=+2c+9 We move all terms containing c to the left and all other terms to the right. +3c-2c=+9-12 We simplify left and right side of the equation. +1c=-3 We divide both sides of the equation by 1 to get c. c=-3

OpenStudy (gelow):

thx

OpenStudy (lizz123):

no prob. need more help tag me